OpenStudy (anonymous):

FOR FUTURE FLVS STUDENTS: (answers are (in order) a, c, b, b, a Question 1 (Worth 1 points) Find the first six terms of the sequence. a1 = -4, an = an-1 + 7 -4, 3, 10, 17, 24, 31 -4, 7, 14, 21, 28, 35 3, 10, 17, 24, 31, 38 0, 7, 14, 21, 28, 35 Points earned on this question: 1 Question 2 (Worth 2 points) Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ... an = -17 + 5(n + 2) an = -17 + 5(n + 1) an = -17 + 5(n - 1) an = -17 x 5(n - 1) Points earned on this question: 2

OpenStudy (anonymous):

Question 3 (Worth 2 points) Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164 an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1) Points earned on this question: 2 Question 4 (Worth 2 points) Find an equation for the nth term of the arithmetic sequence. -1, 2, 5, 8, ... an = -1 + 3(n) an = -1 + 3(n - 1) an = -1 + 3 an = -1 + 3(n + 1) Points earned on this question: 2 Question 5 (Worth 1 points) Find the first six terms of the sequence. a1 = -8, an = 5 • an-1 -8, -40, -200, -1000, -5000, -25,000 -8, -40, -35, -30, -25, -20 0, 5, -40, -35, -30, -25 -40, -200, -1000, -5000, -25,000, -125,000 Points earned on this question: 1 Question 6 (Worth 2 points) A certain species of tree grows an average of 0.5 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 800 centimeters tall. Essay Submission a = 800 cm common difference = 0.5 = 200 + ( n - 1 ) 0.5 = 200 + 0.5n - 0.5 = 199.5 +0.5n